Find the distance between M(2, -1) and N(-4, 3) to the nearest tenth.
Answer by checkley77(12844)

THE X DISTANCE=4+2=6
THE Y DISTANCE=3+1=4
4^2+6^2=C^2
16+36=C^2
C^2=52
C=SQRT52
C=7.2 IS THE DISTANCE BETWEEN THESE 2 POINTS.